If 1.250 of AU and 1.744 of CL2 were mixed determine which of the reactants is in excess
2. Calculate the excess amount
1
Expert's answer
2020-03-30T10:49:41-0400

"2Au + 3Cl_2 \rightarrow 2AuCl_3"

1) Find moles of Au

"1.250 g Au (\frac{1 mole Au}{196.97 g Au})=0.006346 mol"


2) Find moles of Cl2

"1.744 g Cl_2 (\frac{1 mol Cl_2}{2\times 35.45 g Cl_2})=0.02460 mol"


3) Compere the values:


"\frac{moles Au}{2}" and "\frac{moles Cl_2}{3}"


"\frac{0.006346 mol Au}{2}" and "\frac{0.02460 mol Cl_2}{3}"


"0.003173 mol Au < 0.008200 mol Cl_2"

Au - is a limiting reactant

Cl_2 is in excess.


3) Use moles of Au to calculate the moles of Cl_2 consumed in the reaction

"0.006346 mol Au (\frac{3 mol Cl_2}{2 mol Au}) = 0.009519 mol Cl_2"


Find the moles of Cl_2 left after the reaction (the excess amount):

"moles Cl_2 = 0.02460 -0.009519 = 0.01508 mol Cl_2"


Find the mass of "Cl_2" left after the reaction:

"0.01508 mol Cl_2 (\frac{2\times 35.45 g Cl_2}{1 mol Cl_2}) =1.069 g Cl_2"


Answer: excess amount: "0.01508 moles Cl_2" , "1.069 g Cl_2"